Tom took a trip of 1,020 miles . He traveled by train at 55 miles an hour and the same number of hours by plane at 285 mph . How many hours did the trip take?
Answer by malakumar_kos@yahoo.com(315)


Tom took a trip of 1,020 miles . He traveled by train at 55 miles an hour and the same number of hours by plane at 285 mph . How many hours did the trip take
Distance travelled by Tom in train in 1 hr = 55 miles
Distance travelled by Tom by plane in 1 hr = 285 miles
Total distance travelled by Tom in 1 hr = 285+55 = 340 miles
Time taken by Tom to travel a distance of 340 miles = 1 hr
Time taken by Tom to travel a distance of 1,020 miles = 1,020/340 = 3 hrs